# Infinity-MM: Scaling Multimodal Performance with Large-Scale and High-Quality Instruction Data

### v1

@@ Appendix C Data composition
Emu2 (Sun et al., 2024b) 10M Caption
LVIS-Instruct(Gupta et al., 2019) 223K General
LLaVA-CC3M-Pretrain-595K(Li et al., 2024b) 595K General
Visdial(Das et al., 2017) 116K General
Sharegpt4(Chen et al., 2023) 3.2M General
STVQA(Agrawal et al., 2024) 43K General
MMC-INST(Liu et al., 2024a) 500K Doc/Chart/Screen
MathV360K(Shi et al., 2024) 338K Math/Reasoning
MMC-Alignment(Liu et al., 2024a) 250K Doc/Chart/Screen
DocReason(Ye et al., 2024) 26K Doc/Chart/Screen
ALLaVA(Chen et al., 2024a) 1.7M General
Cocotext(Havard et al., 2017) 163K General
Docvqa(Ye et al., 2024) 16K Doc/Chart/Screen
Geoqa+(Chen et al., 2021) 72K Math/Reasoning
DocDownstream(Ye et al., 2024) 700K Doc/Chart/Screen
Cambrian Tong et al. (2024) 8.3M
@@
DocStruct4M(Ye et al., 2024) 4M General OCR, Doc/Chart/Screen
LLaVA-onevision Li et al. (2024a) 4M
@@
Docmatix(Laurençon et al., 2024) 1.2M Doc VQA
Infinity-Instruct BAAI (2024b) 7M Text Instruct
@@
Appendix D Instruction Label System
D.1 Coarse Perception
@@
D.2 Fine-grained Perception (single-instance)
@@
D.3 Fine-grained Perception (cross-instance)

### v2

@@ Appendix B Data composition
Emu2 [59] 10M Caption
LVIS-Instruct[25] 223K General
LLaVA-CC3M-Pretrain-595K[34] 595K General
Visdial[17] 116K General
Sharegpt4[12] 3.2M General
STVQA[2] 43K General
MMC-INST[38] 500K Doc/Chart/Screen
MathV360K[55] 338K Math/Reasoning
MMC-Alignment[38] 250K Doc/Chart/Screen
DocReason[71] 26K Doc/Chart/Screen
ALLaVA[10] 1.7M General
Cocotext[26] 163K General
Docvqa[49] 16K Doc/Chart/Screen
Geoqa+[11] 72K Math/Reasoning
DocDownstream[71] 700K Doc/Chart/Screen
Cambrian [60] 8.3M
@@
DocStruct4M[66] 4M General OCR, Doc/Chart/Screen
LLaVA-onevision [33] 4M
@@
Docmatix[31] 1.2M Doc VQA
Infinity-Instruct [5] 7M Text Instruct
@@
Appendix C Instruction Tagging System
C.1 Coarse Perception
@@
C.2 Fine-grained Perception (single-instance)
@@
C.3 Fine-grained Perception (cross-instance)